Although women make up the majority of nurses, male nurses on average make more than women do. A 2013 report showed that in hospitals, male nurses are bringing in more money per year than female nurses and this disparity increases when we consider ambulatory/outpatient care. If nurses salaries are normally distributed, and a female's salary has a mean of $55,786 with a standard deviation of $4600 and a males salary is $59,686 with a standard deviation of $5100 answer the following:

c. If you are a female nurse, what is the probability that your salary is less than $56,000?

d. If you are a male nurse, what is the probability that your salary is less than $60,000?
Answer #1

a)

for normal distribution z score =(X-μ)/σx
here mean=       μ= 55786
std deviation   =σ= 4600.000
probability =P(X<56000)=(Z<(56000-55786)/4600)=P(Z<0.05)=0.5199

b)

here mean=       μ= 59686
std deviation   =σ= 5100.000
probability =P(X<60000)=(Z<(60000-59686)/5100)=P(Z<0.06)=0.5239